OpenStudy (anonymous):

Someone PLEASE HELP ME Help me simplify the expression below 54/9 * ( 5-2) +32 +8

OpenStudy (anonymous):

please help

OpenStudy (pfenn1):

Take each section of the expression and do the math. so 54/9=6, (5-2)=3, 32+8=40, etc.

OpenStudy (anonymous):

no your supposed to use order of operations

OpenStudy (pfenn1):

I am reading the equation as (54/9) * (5-2) + (32+8) Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Okay, then if you solve each section of the expression you should get (6) * (3) + (40) = ?

OpenStudy (pfenn1):

Does that make sense?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but your suppose to solve the parentheses first

OpenStudy (pfenn1):

I did. (54/9) = 6 (5-2) = 3 (32 + 8) = 40 Am I missing something?